### v1

# The Stepanov theorem for Q-valued functions

Abstract.
In this work we prove the Stepanov differentiation theorem for multiple-valued functions. This theorem is proved in the wide generality of metric-space-multiple-valued functions without relying on a Lipschitz extension result. General definitions of differentiability and approximate differentiability for functions between suitable metric spaces are also introduced.
Key words and phrases:
Lipschitz functions, Rademacher’s Theorem, Multiple-valued functions, Stepanov’s theorem, Metric measure spaces.
2020 Mathematics Subject Classification:
54E40, 54E45, 46E36, 26B05, 26B35.
The author is supported by the University of Padova, in particular through the STARS@unipd project “QuASAR - Questions About Structure And Regularity of currents” (MASS_STARS_MUR22_01).
area-minimizing
1. Introduction
$Q$-valued functions has been developed by Almgren to study regularity properties of mass-minimizing currents. In particular, in [1] he generalized the concepts of differential and Sobolev space for $Q$-valued functions, i.e. functions with values in the space $\mathcal{A}_{Q}\left(M\right)$ made by unordered $Q$-tuple of points in a metric space $M$, in order to study regularity properties of “Dir-minimizing” functions, which are $Q$-valued functions that minimize a suitable generalization of the Dirichlet energy for $Q$-valued functions. His work about Dir-minimizing functions has been readapted and simplified in [3].
When $M={\mathbb{R}}^{k}$, we say that a $Q$-valued function ${f}:{\Omega\subseteq{\mathbb{R}}^{n}}\rightarrow{\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)}$ is differentiable at $x\in\Omega$ with $f(x)=\left\llbracket P_{1}\right\rrbracket+\left\llbracket P_{2}\right\rrbracket+\dotsb+\left\llbracket P_{Q}\right\rrbracket$ if and only if there exist $M_{1},M_{2},\dotsc,M_{Q}$ $(k\times n)$-matrix such that, setting
$F(z)=\sum^{Q}_{i=1}\left\llbracket P_{i}+M_{i}z\right\rrbracket\text{,}$
we have
$\lim_{y\to x}\frac{\mathcal{G}\left[f(y),F(y-x)\right]}{\left\lvert y-x\right\rvert}=0$
(where $\mathcal{G}\left[\cdot,\cdot\right]$ is the canonical metric on $\mathcal{A}_{Q}\left(M\right)$ as defined in section 3) and if $P_{i}=P_{j}$ for some $i,j$ then $M_{i}=M_{j}$.
In [3] the following version of the Rademacher theorem for $Q$-valued functions is proved:
Theorem 1.1.
If ${f}:{{\mathbb{R}}^{n}}\rightarrow{\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)}$ is a Lipschitz $Q$-valued function then $f$ is a.e. differentiable on ${\mathbb{R}}^{n}$.
To prove it the authors used a Lipschitz extension result for $Q$-valued Lipschitz functions from a closed subset of ${\mathbb{R}}^{n}$ to all ${\mathbb{R}}^{n}$. However, we cannot expect a Lipschitz extension result when we replace ${\mathbb{R}}^{k}$ with a generic metric space, because it is not always possible to globally extend a continuous function defined on a closed subset of ${\mathbb{R}}^{n}$ with values on a disconnected metric space.
However, an extension result for $Q$-valued Lipschitz functions is not strictly required, since it is possible to prove theorem 1.1 without using the extension result for Lipschitz $Q$-valued functions at the cost of introducing a slightly weaker notion of differentiable function (which coincides with the classical one on interior points).
The main goal of this work is to prove the the Stepanov theorem for $Q$-valued functions:
Theorem 1.2.
Let ${f}:{B}\rightarrow{\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)}$ be a measurable function defined on a measurable set $B\subseteq{\mathbb{R}}^{n}$ without isolated points. Let
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{y\in B\to x}\frac{\mathcal{G}\left[f(y),f(x)\right]}{\left\lvert y-x\right\rvert}$
$A_{f}=\left\{{x\in B}\mathrel{}\middle|\mathrel{}{\left\lvert\nabla f\right\rvert^{+}(x)<+\infty}\right\}\text{.}$
Then $f$ is a.e. differentiable on $A_{f}$.
However, we will generalize these concepts in order to functions defined on doubling metric measure spaces with values on complete metric spaces and not necessarly measurable.
In Section 2 of this work we introduce a broad definition of “differential classes” for functions defined between two metric spaces. This definition generalizes the usual definition of differentiation for single-valued functions defined between euclidean spaces, $Q$-valued functions and real-valued functions defined on $p$-Poincaré metric measuree spaces (see [2], [5] and [4]). In particular, we say that a function $f$ defined on a nonempty subset of a doubling metric measure space with values on a complete metric space is differentiable at some accumulation point $x$ if and only if its first-order asymptotical behaviour near $x$ belongs to a suitable “differential structure” that makes every Lipschitz function defined on closed subsets differentiable almost everywhere. This Rademacher-type condition is all we need to prove Stepanov’s theorem for functions between metric spaces.
Section 3 is instead entirely devoted to differentiability properties of metric $Q$-valued functions, that is functions with values in $\mathcal{A}_{Q}\left(M\right)$ instead of $\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)$ with $M$ any complete metric space, and to find sufficient conditions for the Rademacher-type condition stated in previous section in order to prove a more general version of theorem 1.2.
2. Generalized differentiability
2.1. Differential germs
Let $(X,\mathtt{J},\mu)$ be a $K$-doubling metric measure space ($K>0$), that is, $(X,\mathtt{J})$ is a complete separable metric space and ${\mu}:{\mathcal{P}(X)}\rightarrow{\left[{0},{+\infty}\right]}$ is a Borel-regular outer measure on $X$ such that for every $x\in X$ and every $r>0$ we have
$0<\mu\left[B_{2r}(x)\right]\leq K\mu\left[B_{r}(x)\right]<+\infty\text{.}$
Remark.
It is simple to show that for every $x,y\in X$ and $r,s>0$
$\mu\left[B_{r}(x)\right]\leq\mathcal{L}\left(K,\frac{r+\mathtt{J}(x,y)}{s}\right)\mu\left[B_{s}(y)\right]$
where $\mathcal{L}(\alpha,\beta)>0$ is a suitable funcion. Indeed we have
$\mu\left[B_{r}(x)\right]\leq\mu\left[B_{r+\mathtt{J}(x,y)}(y)\right]\leq\mu\left[B_{2^{P\left[\left(r+\mathtt{J}(x,y)\right)/s\right]}s}(y)\right]\\ \leq K^{P\left[\left(r+\mathtt{J}(x,y)\right)/s\right]}\mu\left[B_{s}(y)\right]$
where $P(t)=\max\left\{{0,\lceil\log_{2}t\rceil}\right\}$.
Proposition 2.1.
Let $E\subseteq X$ be any set, there exists a Borel sets $B\supseteq E$ such that $\mu\left[B^{\prime}\cap B_{r}(x)\right]\geq\mu\left[B\cap B_{r}(x)\right]$ for every $x\in X,r>0$ and every other measurable set $B^{\prime}\supseteq E\cap B_{r}(x)$.
Proof.
Fix a point $x_{0}\in X$ and set $U_{r}=B_{r}\left(x_{0}\right)$ for every $r>0$. By hypothesis for every $r>0$ we have $0<\mu\left(U_{r}\right)<+\infty$ and there exists $E\cap U_{r}\subseteq B_{r}\subseteq U_{r}$ Borel set such that $\mu\left(B_{r}\right)=\mu\left(E\cap U_{r}\right)$ since $\mu$ is Borel-regular. Then for every $B_{s}(y)\subseteq U_{r}$ we have
$\begin{split}\mu\left(E\cap U_{r}\right)&\leq\mu\left(E\cap B_{s}(y)\right)+\mu\left(E\cap U_{r}\setminus B_{s}(y)\right)\\ &\leq\mu\left(B_{r}\cap B_{s}(y)\right)+\mu\left(B_{r}\setminus B_{s}(y)\right)\\ &=\mu\left(B_{r}\right)=\mu\left(E\cap U_{r}\right)\end{split}$
which implies $\mu\left(E\cap B_{s}(y)\right)=\mu\left(B_{r}\cap B_{s}(y)\right)$ for every $B_{s}(y)\subseteq U_{r}$.
Now let $B=\bigcup^{\infty}_{S=1}\bigcap^{\infty}_{R=S}B_{R}$, clearly $E\subseteq B$ and $B$ is a Borel set. For every ball $B_{s}(y)$ there exists $\tilde{S}\in{\mathbb{N}}$ such that $B_{s}(y)\subseteq U_{\tilde{S}}$ and so
$\mu\left(B\cap B_{s}(y)\right)=\lim_{S\to+\infty}\mu\left(\bigcap^{+\infty}_{R=S}\left(B_{R}\cap B_{s}(y)\right)\right)\leq\mu\left(E\cap B_{s}(y)\right)\text{.}$
Therefore $\mu\left(B\cap B_{s}(y)\right)=\mu\left(E\cap B_{s}(y)\right)$. ∎
A point $x$ of $E$ is said to be a $\mu$-interior point (for $E$) if and only if there exists $B\subseteq E$ Borel set such that
$\lim_{r\to 0^{+}}\frac{\mu\left[B_{r}(x)\setminus B\right]}{\mu\left[B_{r}(x)\right]}=0\text{.}$
Let also ${E}^{1}\subseteq E$ be the set of all the $\mu$-interior points for $E$, and if $x\in{E}^{1}$ then we say that $E$ is a $\mu$-neighborhood for $x$.
Notice that $E$ and ${E}^{1}$ may not be $\mu$-measurable sets, however if $E$ is $\mu$-measurable then $\mu\left(E\setminus{E}^{1}\right)=0$ by using the doubling property of $\mu$ (see for example [6]). Moreover, if $E$ and $F$ are two $\mu$-neighborhoods for $x$ then $E\cap F$ is another $\mu$-neighborhood for $x$ because
$\lim_{r\to 0^{+}}\frac{\mu\left[B_{r}(x)\setminus(E\cap F)\right]}{\mu\left[B_{r}(x)\right]}\leq\lim_{r\to 0^{+}}\frac{\mu\left[B_{r}(x)\setminus E\right]}{\mu\left[B_{r}(x)\right]}+\lim_{r\to 0^{+}}\frac{\mu\left[B_{r}(x)\setminus F\right]}{\mu\left[B_{r}(x)\right]}=0\text{.}$
Therefore the family of $\mu$-neighborhoods defines a topology on $X$, that we will call the $\mu$-topology. This topology is clearly stronger than the metric topology induced by $\mathtt{J}$ on $X$ (which we will also call $\mathtt{J}$-topology).
Proposition 2.2.
For every $x\in X$ all the following statements are equivalent:
  • $x$ is an isolated point (for $X$) with respect to the $\mathtt{J}$-topology;
  • $x$ is a $\mu$-isolated point;
  • $\mu\left(\left\{{x}\right\}\right)>0$.
Proof.
If $x$ is an isolated point then there exists $r>0$ such that $B_{r}(x)=\left\{{x}\right\}$, and since $B_{r}(x)$ is itself a $\mu$-neighborhood for $x$ then this would automatically imply that $x$ is an isolated point.
Next if $x$ is $\mu$ isolated then $\left\{{x}\right\}$ is a $\mu$-neighborhood for $x$ so
$\lim_{r\to 0^{+}}\frac{\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]}{\mu\left[B_{r}(x)\right]}=0$
which is possible only if $\mu(\left\{{x}\right\})>0$ otherwise $\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]=\mu\left[B_{r}(x)\right]$.
Finally we only need to prove that if $\mu(\left\{{x}\right\})=\delta>0$ then $x$ is an isolated point, that is $B_{r}(x)=\left\{{x}\right\}$ for some $r>0$. First to all notice that, since $\mu\left[B_{r}(x)\right]<+\infty$, $\lim_{r\to 0^{+}}\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]=0$. Therefore there exists $\tilde{r}>0$ such that $\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]<\delta/K$ for every $0<r<\tilde{r}$.
By contradiction $x$ is not an isolated point in $X$, therefore fixed $0<r<\tilde{r}$ there exist $y\neq x$ and $s>0$ such that
  • $B_{s}(y)\subseteq B_{r}(x)\setminus\left\{{x}\right\}$;
  • $x\in B_{2s}(y)$.
Then
$\mu\left[B_{s}(y)\right]\leq\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]<\frac{\delta}{K}=\frac{1}{K}\mu\left(\left\{{x}\right\}\right)\leq\frac{1}{K}\mu\left[B_{2s}(y)\right]\leq\mu\left[B_{s}(y)\right]$
which is impossible, therefore $\left\{{x}\right\}$ must be an isolated point. ∎
Proposition 2.3.
Consider any nonempty set $U\subseteq X$. Then for every $x\in{U}^{1}$ we can find an application ${r_{x}}:{X}\rightarrow{U}$ such that $r_{x}(y)=y$ on $U$ and
$\mathtt{J}\left[y,r_{x}(y)\right]=o\left(\mathtt{J}\left(y,x\right)\right)\text{.}$
Proof.
We can assume without loss of generality that $x$ is an accumulation point (that is, $B_{r}(x)\neq\left\{{x}\right\}$ for every $r>0$ or equivalently $\mu(\left\{{x}\right\})=0$) and $U$ is a Borel set. For every $N\in{\mathbb{N}}$ we can find $\delta_{N}>0$ such that for every $0<r<\delta_{N}$
$\frac{\mu\left[B_{r}(x)\setminus U\right]}{\mu\left[B_{r}(x)\right]}\leq\frac{1}{N}\text{.}$
In order to prove our statement we need only to prove that for every $\varepsilon>0$ there exists $\delta>0$ so that for every $y\in X$ with $\left\lvert y-x\right\rvert<\delta$ there exists $y^{\prime}\in U$ such that
$\mathtt{J}\left(y,y^{\prime}\right)<\varepsilon\mathtt{J}\left(y,x\right)\text{.}$ (2.1)
If (2.1) does not hold then there exists $\varepsilon_{0}>0$ such that for every $N\in{\mathbb{N}}$ there exists $y_{N}$ such that $0<r_{N}=\mathtt{J}\left(y_{N},x\right)<\delta_{N}/\left(1+\varepsilon_{0}\right)$ but $B_{\varepsilon_{0}r_{N}}\left(y_{N}\right)\cap U=\emptyset$. Therefore, since we have $\left(1+\varepsilon_{0}\right)r_{N}<\delta_{N}$,
$\mu\left[B_{\varepsilon_{0}r_{N}}\left(y_{N}\right)\right]\leq\left\lvert B_{\left(1+\varepsilon_{0}\right)r_{N}}(x)\setminus U\right\rvert\leq\frac{1}{N}\mu\left[B_{(1+\varepsilon_{0})r_{N}}(x)\right]\\ \leq\frac{1}{N}\mu\left[B_{(2+\varepsilon_{0})r_{N}}(y_{N})\right]\leq\frac{K_{\varepsilon_{0}}}{N}\mu\left[B_{\varepsilon_{0}r_{N}}(y_{N})\right]$
where $K_{\varepsilon_{0}}$ is a constant that depends only on $\varepsilon_{0}>0$ and $\mu$ but not on $N$. Therefore, for $N$ sufficiently large we will get a contradiction. ∎
Limits in the $\mu$-topology coincide with the approximate limits with respect to measure $\mu$, that is let ${g}:{A}\rightarrow{{\mathbb{R}}}$ be a function defined on a nonempty set $A\subseteq X$ and let $x\in X$ be a $\mu$-accumulation point for $A$ (that is, $A\cap U\setminus\left\{{x}\right\}\neq\emptyset$ for every $\mu$-neighborhood $U$ of $x$). Then we have $\mu-\lim_{y\in A\to x}g(y)=z\in{\mathbb{R}}$ if and only if for every $\varepsilon>0$ there exists a $\mu$-neighborhood $U$ of $x$ such that $\left\lvert g(y)-z\right\rvert<\varepsilon$ for every $y\in A\cap U\setminus\left\{{x}\right\}$.
Moreover, for real-valued functions ${g}:{A}\rightarrow{{\mathbb{R}}}$ and $\mu$-accumulation point $x$ for $A\subseteq X$ we set
$\mu-\limsup_{y\in A\to x}g(y)=\inf_{x\in{U}^{1}}\sup_{y\in U\cap A\setminus\left\{{x}\right\}}g(y)\text{.}$
Proposition 2.4.
Let ${f}:{B}\rightarrow{\left[{0},{+\infty}\right)}$ be a real-valued Lipschitz functions defined on a $\mu$-neighborhood $B$ of $x$ with $\mu(\left\{{x}\right\})=0$ and $f(x)=0$. Then
$\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}=\mu-\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\text{.}$
Proof.
The inequality $\geq$ is trivial, to prove the opposite inequality take any $\mu$-neighborhood $U$ of $x$ and set $C=U\cap B\setminus\left\{{x}\right\}$. Let ${r_{x}}:{X\setminus\left\{{x}\right\}}\rightarrow{C}$ as in proposition 2.3, then for every $\varepsilon>0$ there exists $\delta>0$ such that if $y\in B_{\delta}(x)$ then $\mathtt{J}[r_{x}(y),y]\leq\varepsilon\mathtt{J}(y,x)$. Therefore,
$\sup_{y\in B_{\delta}(x)\cap B\setminus\left\{{x}\right\}}\frac{f(y)}{\mathtt{J}(y,x)}\leq\sup_{y\in C}\frac{f(y)}{\mathtt{J}(y,x)}\sup_{y\in B_{\delta}(x)\setminus\left\{{x}\right\}}\frac{\mathtt{J}\left[r_{x}(y),x\right]}{\mathtt{J}(y,x)}\\ +\sup_{y\in B_{\delta}(x)\cap B\setminus\left\{{x}\right\}}\frac{\left\lvert f(y)-f[r_{x}(y)]\right\rvert}{\mathtt{J}(y,x)}\leq(1+\varepsilon)\sup_{y\in C}\frac{f(y)}{\mathtt{J}(y,x)}+K\varepsilon$
for some $K>0$. Thus
$\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\leq(1+\varepsilon)\mu-\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}+K\varepsilon$
which proves our statement. ∎
In the same way we also set
$\left\lvert\nabla f\right\rvert^{\mu,+}(x)=\mu-\limsup_{y\in A\to x}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}(x,y)}$
$A_{f}^{\mu}=\left\{{x\in A}\mathrel{}\middle|\mathrel{}{x\ \mu\text{-accumulation point, }\left\lvert\nabla f\right\rvert^{\mu,+}(x)<+\infty}\right\}\text{.}$
Clearly we have $\left\lvert\nabla f\right\rvert^{\mu,+}(x)\leq\left\lvert\nabla f\right\rvert^{+}(x)$ and $A_{f}\subseteq A_{f}^{\mu}\cup N$ where $\mu(N)=0$, because $A_{f}\setminus A_{f}^{\mu}$ may contain accumulation points for $A$ that are $\mu$-isolated. This set should have measure zero since $\mu$ is Borel regular and so there exists $\tilde{A}\supseteq A$ Borel set such that $\mu\left[B_{r}(x)\cap\tilde{A}\right]=\mu\left[B_{r}(x)\cap A\right]$ for every $x\in X$, $r>0$.
Even if $f$ is a Lipschitz function defined on a closed set $C\subseteq X$ we can have $A_{f}^{\mu}\neq C$, in this case $C\setminus A_{f}^{\mu}$ is equal to all the $\mu$-isolated points of $C$.
Take now another complete metric space $\left(M,d\right)$, we want to define a notion of differentiability for functions defined on $X$ with values on $M$. For every topology $\tau$ on $X$ stronger than or equal to the $\mathtt{J}$-topology (in particular, we set $\tau=\mathtt{J}$ for the $\mathtt{J}$-topology and $\tau=\mu$ for the $\mu$-topology), every $x\in X$, $P\in M$ let $\mathrm{Lip}^{\tau}_{x,P}\left(M,X\right)$ be the set of all the equivalence class $\left[F\right]_{\sim}$, where ${F}:{B}\rightarrow{M}$ is a Lipschitz function defined on a $\tau$-neighborhood $B\subseteq X$ of $x$ such that $F(x)=P$, with respect to the equivalence relation
$F\sim G\Leftrightarrow\lim_{y\to x}\frac{d[F(y),G(y)]}{\mathtt{J}\left(y,x\right)}=0\text{ (with respect to }\tau\text{ topology).}$).
For every $(x,P)\in X\times M$ the space $\left[\mathrm{Lip}^{\tau}_{x,P}\left(M,X\right),d_{x,P},0_{P},\left\lvert\cdot\right\rvert_{x,P}\right]$ is a pointed metric space where
$d_{x,P}([f],[g])=\limsup_{y\to 0}\frac{d[f(y),g(y)]}{\mathtt{J}\left(y,x\right)}$
$0_{P}=\left[{x\in X}\rightarrow{P\in M}\right]$
$\left\lvert[f]\right\rvert_{x,P}=d_{x,P}\left([f],0_{P}\right)\text{.}$
We also define
$\mathrm{Lip}^{\tau}\left(M,X\right)=\bigsqcup_{(x,P)\in X\times M}\mathrm{Lip}^{\tau}_{x,P}\left(M,X\right)$
and we call the elements of $\mathrm{Lip}^{\tau}\left(M,X\right)$ the $X$-differential germs of $M$ with respect to the topology $\tau$. In particular, if $\tau=\mathtt{J}$ then the elements of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are just called $X$-differential germs, whereas if $\tau=\mu$ then elements of $\mathrm{Lip}^{\mu}\left(M,X\right)$ are called approximate $X$-differential serms.
When $\tau$ is the topology induced by the metric $\mathtt{J}$ then arbitrary subsets $\mathcal{D}$ of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are called $X$-distributions for $M$, and for every $(x,P)\in X\times M$ we set $\mathcal{D}_{x,P}=\mathcal{D}\cap\mathrm{Lip}^{\mathtt{J}}_{x,P}\left(M,X\right)$. Analogously, when $\tau$ is the topology generated by the measure $\mu$ as before then arbitrary subsets of
$\mathrm{Lip}^{\mu}\left(M,X\right)$ are instead called approximate $X$-distributions.
2.2. Differentiaition and completeness
Definition 2.5.
Let $A$ be a nonempty subset of $X$, $\mathcal{D}$ be an $X$-distribution and let ${f}:{A}\rightarrow{M}$ be an arbitrary function. For every $x\in A$ accumulation point we say that $f$ is $\mathcal{D}$-differentiable at $x$ if and only if there exists $[{F}:{B}\rightarrow{M}]\in\mathcal{D}_{x,f(x)}$ such that
$\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0\text{.}$ (2.4)
Moreover, the differential germ $[F]$ that satisfies (2.4) is a $\mathcal{D}$-differential of $f$ at $x$.
If $\mathcal{D}$ is an approximate $X$-distribution and $x$ is a $\mu$-accumulation point for $A$ then $f$ is approximately $\mathcal{D}$-differentiable at $x$ if and only if
$\mu-\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}=0$
and in that case $[F]$ is an approximate $\mathcal{D}$-differential for $f$ at $x$.
Remark.
If $x$ is a $\mu$-interior point for $A$ then $f$ admits at most one approximate differential at $x$. However, if $x$ is a boundary point then the differential may not be unique and $f$ could have infinitely many different differentials.
Proposition 2.6.
Let ${f}:{A}\rightarrow{M}$ be an arbitrary function that is $\mathcal{D}$-differentiable at some accumulation point $x\in A$. Then $x\in A_{f}$ and $\left\lvert\nabla f\right\rvert^{+}(x)\leq\left\lvert F\right\rvert_{x,f(x)}$ where $[F]\in\mathcal{D}_{x,f(x)}$ is any differential for $f$.
If in addition $A$ is a $\mu$-neighborhood for $x$ then $\left\lvert\nabla f\right\rvert^{+}(x)=\left\lvert T_{x}f\right\rvert_{x,f(x)}$ where $T_{x}f=\left[F\right]\in\mathcal{D}_{x,f(x)}$ is the unique differential of $f$ at $x$.
Proof.
Let $[F]\in\mathcal{D}_{x,f(x)}$ be a $\mathcal{D}$-differential of $f$ at $x$, then
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}\left(y,x\right)}\\ \leq\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[f(y),F(y)\right]}{\mathtt{J}\left(y,x\right)}+\limsup_{y\to x}\frac{d\left[F(y),F(x)\right]}{\mathtt{J}\left(y,x\right)}=\left\lvert F\right\rvert_{x,f(x)}\text{.}$
Next we prove that if $x$ is a $\mu$-interior point for $A$ then $f$ has exactly one $\mathcal{D}$-differential at $x$ and equality holds in the last inequality. Let $[G]\in\mathcal{D}_{x,f(x)}$ be another $\mathcal{D}$-differential of $f$ at $x$, then we immediately get
$\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0$
but this does not yet imply that $[F]=[G]$. Since $A$ is a $\mu$-neighborhood of $x$ we immediately get
$\mu-\limsup_{y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=\mu-\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}\\ \leq\lim_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0\text{.}$
Now both $F$ and $G$ are defined on a neighborhood of $x$, by proposition 2.4 we finally get
$\limsup_{y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=\mu-\limsup_{y\to x}\frac{d\left[F(y),G(y)\right]}{\mathtt{J}(y,x)}=0$
thus $[F]=[G]$.
Assime again that $A$ is a $\mu$-neighborhood for $x$, and let $T_{x}f=[F]$ be the unique $\mathcal{D}$-differential of $f$ at $x$. Then
$\left\lvert T_{x}f\right\rvert_{x,f(x)}=\mu-\limsup_{y\to x}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\\ =\mu-\limsup_{\begin{subarray}{c}y\to x\\ y\in A\end{subarray}}\frac{d[F(y),F(x)]}{\mathtt{J}(y,x)}\leq\left\lvert\nabla f\right\rvert^{+}(x)\text{.}\qed$
In a similar fashion we can prove the following:
Proposition 2.7.
Let ${f}:{A}\rightarrow{M}$ be an arbitrary function that is approximately $\mathcal{D}$-differentiable at some $x\in A$. Then $x\in A_{f}^{\mu}$ and $\left\lvert\nabla f\right\rvert^{\mu,+}(x)\leq\left\lvert F\right\rvert_{x,f(x)}$ where $[F]\in\mathcal{D}_{x,f(x)}$ is any approximate differential for $f$.
If in addition $A$ is a $\mu$-neighborhood for $x$ then the approximate $\mathcal{D}$-differential $T_{x}f=[F]$ of $f$ at $x$ is unique with $\left\lvert\nabla f\right\rvert^{\mu,+}(x)=\left\lvert T_{x}f\right\rvert_{x,f(x)}$.
We then say that a function ${f}:{A}\rightarrow{M}$ defined on a nonempty set $A\subseteq X$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$ if and only if on every $\mu$-measurable subset $B\subseteq A_{f}$ $\mu$-measurable we have that $f$ is $\mathcal{D}$-differentiable for a.e. point of $B$. In the same way we define functions a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
Definition 2.8.
An $X$-distribution/approximate $X$-distribution $\mathcal{D}$ of $M$ is differentiably complete if it satisfies the following “Rademacher’s condition”:
(RC):
for every closed nonempty subset $C\subseteq X$ (with respect to the $\mathtt{J}$-topology) and every Lipschitz function ${g}:{C}\rightarrow{M}$ we have that $g$ is a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
Remark.
For Lipschitz functions $\mathcal{D}$-differentiability and approximate $\mathcal{D}$-differentiability are equivalent for a.e. point of $C$, where $\mathcal{D}$ is an $X$-distribution. Moreover, both $A_{f}$ and $A_{f}^{\mu}$ are Borel sets with $\mu\left({A_{f}}\Delta{A_{f}^{\mu}}\right)$ $=0$.
Remark.
In (RC) we cannot assume $C=X$ since for a generic metric space $M$ we do not have a local Lipschitz extension result.
It is trivial to show that
  1. (1)
    $\mathcal{D}=\mathrm{Lip}^{\mu}\left(M,X\right)$ is differentiably complete;
  2. (2)
    if $\mathcal{D}$ is differentiably complete then every $\mathcal{D}\subseteq\mathcal{C}$ is also differentiably complete;
  3. (3)
    if $\mathcal{D}^{1},\mathcal{D}^{2},\mathcal{D}^{3},\dotsc$ are all differentiably complete then $\mathcal{D}=\bigcap^{\infty}_{i=1}\mathcal{D}^{i}$ is differentiably complete.
Of these statements, only point 3 is not trivial, however its proof is quite simple thanks to uniqueness results contained in propositions 2.6 and 2.7. Let then ${f}:{C}\rightarrow{M}$ be a Lipschitz function defined on a closed subset $C$ of $X$, then there exists a measurable subset $V={V}^{1}$ of $C$ such that for every $i\in{\mathbb{N}}$ $f$ is approximately $\mathcal{D}^{i}$-differentiable at $x$ with differential $\left[F_{i}\right]\in\mathcal{D}^{i}$. By uniqueness we have $\left[F_{i}\right]=\left[F_{1}\right]=T_{x}f$ for every $i\in{\mathbb{N}}$, therefore $T_{x}f\in\bigcap^{\infty}_{i=1}\mathcal{D}^{i}_{x,f(x)}=\mathcal{D}_{x,f(x)}$.
Example 2.9.
Let $\left\{{X_{\alpha}}\right\}_{\alpha\in{\mathbb{N}}}$ be a disjoint Borel cover of $(X,\mathtt{J},\mu)$ and let ${\Phi_{\alpha}}:{X_{\alpha}}\rightarrow{{\mathbb{R}}^{N_{\alpha}}}$ be a sequence of Lipschitz functions with $N_{\alpha}\in{\mathbb{N}}$.
We also assume that $\mu\left(X_{\alpha}\right)>0$ and that $X_{\alpha}$ does not contain any $\mu$-isolated point (with respect to itself). Therefore for every $k\in{\mathbb{N}}$, every $x\in{\left(X_{\alpha}\right)}^{1}$ and every $p\in{\mathbb{R}}^{k}$ we can set
$\mathcal{D}_{x,p}=\left\{{\left[{y\in X_{\alpha}}\rightarrow{p+L\left[\Phi_{\alpha}(y)-\Phi_{\alpha}(x)\right]}\right]}\mathrel{}\middle|\mathrel{}{{L}:{{\mathbb{R}}^{N_{\alpha}}}\rightarrow{{\mathbb{R}}^{k}}\text{ linear}}\right\}\text{.}$
Clearly $\mathcal{D}=\bigsqcup_{(x,p)\in X\times{\mathbb{R}}^{k}}\mathcal{D}_{x,p}$ is an approximate $X$-distribution for ${\mathbb{R}}^{k}$ and an $X_{\alpha}$-distribution for ${\mathbb{R}}^{k}$ for every $\alpha$.
A well-know result proven for the first time by Cheeger ([2]) states that under some conditions on the base space $X$ (in particular, the existence of a $p$-Poincaré type inequality for some $1\leq p<+\infty$) then $\mathcal{D}$ is a complete distribution for every $k\in{\mathbb{N}}$ and $\sup N_{\alpha}<+\infty$, in other words the metric space $X$ is “locally finite dimensional”.
In the last part of this section we will prove the equivalence of (RC) with a $\mathcal{D}$-version of the classical Stepanov theorem for differentiability and approximate differentiability. We first need to generalize proposition 2.3.
Let $\delta,\eta>0$, we say a measurable set $B\subseteq X$ containing $y$ is $(\delta,\eta)$-full at $y$ if
$\sup_{0<r<\eta}\frac{\mu\left[B_{r}(y)\setminus B\right]}{\mu\left[B_{r}(y)\right]}\leq\delta\text{.}$
Lemma 2.10.
There exists $\delta=\delta(K)>0$ such that the following statements hold:
  1. (1)
    if $B_{x}$ and $B_{y}$ are $(\delta,\eta)$-full for some $\eta>0$ at $x$ and $y$ respectively and $30<t=\mathtt{J}(x,y)<\eta/$ 3 then $\mu\left[B_{3t}(x)\cap B_{x}\cap B_{3t}(y)\cap B_{y}\right]>0$;
  2. (2)
    let $E\subseteq X$ be a nonempty set such that $x\in{\left(\overline{E}\right)}^{1}$ and for every $y\in E$ let $B_{y}$ be a measurable set so that for every $z\in\overline{E}\setminus\left\{{x}\right\}$ sufficiently close to $x$ there exist $l,\eta>0$ such that $B_{y}$ is $(\delta,\eta)$-full at $y$ for every $y\in B_{l}(z)$, then there exist a $\mu$-neighborhood $U$ of $x$ and an application ${r_{x}}:{U}\rightarrow{E}$ such that $y\in B_{r_{x}(y)}$ for every $y\in U\setminus\left\{{x}\right\}$.
Proof.
Fix $\delta>0$ and assume by contradiction that 1. does not hold. Thus $\mu\left[B_{y}\cap B_{t}(y)\right]\leq\mu\left[B_{3t}(y)\setminus B_{y}\right]$ with $3t<\eta$ and then
$\begin{split}(1-\delta)\mu\left[B_{t}(x)\right]&\leq\mu\left[B_{t}(x)\cap B_{x}\right]\\ &\leq\mu\left[B_{3t}(y)\setminus B_{y}\right]\\ &\leq\delta\mu\left[B_{3t}(y)\right]\\ &\leq\delta\mathcal{L}(K,4)\mu\left[B_{t}(x)\right]\end{split}$
which is impossible when $\delta<1/\left[1+\mathcal{L}(K,4)\right]$.
To prove 2 we can assume without loss of generality that $x$ is a $\mu$-accumulation point, that is $\mu(\left\{{x}\right\})=0$. Notice also that we can find a sequence $y_{n}\in E$ such that $\overline{E}=\overline{\left\{{y_{n}}\mathrel{}\middle|\mathrel{}{n\in{\mathbb{N}}}\right\}}$. Then we need only to prove that
$\mu\left[\overline{E}\cap B_{\xi}(x)\setminus\left(\left\{{x}\right\}\cup\bigcup_{n\in{\mathbb{N}}}B_{y_{n}}\right)\right]=0$
for $\xi>0$ sufficiently small.
Otherwise we can find $Z\subseteq\overline{E}$ sufficiently near $x$ such that $Z\cap B_{y_{n}}=\emptyset$ and
$\sup_{0<r<\eta_{0}}\frac{\mu\left[B_{r}(z)\cap Z\right]}{\mu\left[B_{r}(z)\right]}\geq 1-\delta$
for some $z\in Z\setminus\left\{{x}\right\}$ and $0<\eta_{0}$, in particular $Z$ if $(\delta,\eta_{0})$-full at $z$. By assumption we can find $l>0$ and $0<\eta<\eta_{0}$ such that $B_{y_{n}}$ is $(\delta,\eta)$-full at $y_{n}$ for every $y_{n}\in B_{l}(z)$. Moreover, for $n$ large we also have $\mathtt{J}\left(y_{n},z\right)<\eta/3$ therefore by 1. we would have $Z\cap B_{y_{n}}\neq\emptyset$, a contradiction. ∎
Theorem 2.11.
Let $\mathcal{D}$ be a differentiably complete approximate $X$-distribution for $M$, then every function ${f}:{A}\rightarrow{M}$ is a.e. approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$.
Conversely let $\mathcal{D}$ be an approximate $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is approximately $\mathcal{D}$-differentiable on $A_{f}^{\mu}$, then $\mathcal{D}$ is differentiably complete.
We first prove that if $\mathcal{D}$ is complete then ${f}:{A}\rightarrow{M}$ is a.e. approximately $\mathcal{D}$-differentiable on every measurable subset $B$ of $\mathcal{A}_{f}^{\mu}$. Fix a point $P\in M$ and for every $i,j\in{\mathbb{N}}$ let $E_{ij}\subseteq A$ such that $x\in E_{ij}$ if and only if
  1. (1)
    $d[f(x),P]<\frac{i}{2j}$;
  2. (2)
    there exists a $(\delta,3/j)$-full set $B_{x}\subseteq A$ at $x$ such that $d[f(y),f(x)]$ $\leq i\mathtt{J}(y,x)$ for every $y\in B_{x}$.
The main difficulty here is that the sets $E_{ij}$ are not in general measurable. By definition $A_{f}^{\mu}=\bigcup_{ij}E_{ij}$, we now prove that ${f_{ij}}:{x\in E_{ij}}\rightarrow{f(x)\in M}$ is $6i$-Lipschitz.
Take any $y,z\in E_{ij}$, if $\mathtt{J}\left(y,z\right)\geq 1/j$ then
$d[f(y),f(z)]\leq d[f(y),P]+d[P,f(z)]\leq\frac{i}{j}\leq i\mathtt{J}\left(y,z\right)\text{.}$
If instead $0<t=\mathtt{J}(y,z)<1/j$ then by lemma 2.10 there exists $s\in B_{z}\cap B_{y}\cap B_{3t}(z)\cap B_{3t}(y)$ such that
$d[f(y),f(z)]\leq d[f(y),f(s)]+d[f(s),f(z)]\\ \leq i\mathtt{J}(y,s)+i\mathtt{J}(s,z)<6i\mathtt{J}(y,z)\text{.}$
Since $M$ is complete we can uniquely extend $f_{ij}$ on $\overline{E_{ij}}$, therefore there exists a measurable subset $\tilde{G}_{ij}$ of $\overline{E_{ij}}$ such that $\mu\left(\overline{E_{ij}}\setminus\tilde{G}_{ij}\right)=0$ and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on each point of $\tilde{G}_{ij}$.
Let $B_{ij}=\overline{E_{ij}}\cap B$ and $G_{ij}={\left(\tilde{G}_{ij}\cap B\right)}^{1}$, then we immediately get
$\mu\left(B_{ij}\setminus G_{ij}\right)\leq\mu\left(\overline{E_{ij}}\setminus\tilde{G}_{ij}\right)+\mu\left(\tilde{G}_{ij}\setminus{\tilde{G}_{ij}}^{1}\right)=0$
and $f_{ij}$ is approximately $\mathcal{D}$-differentiable on every point of $G_{ij}$. Let then $x\in G_{ij}$ and $\left[F_{ij}\right]\in\mathcal{D}_{x,f(x)}$ be the differential of $f_{ij}$ in $x$, then for every $\varepsilon>0$ there exists a measurable $\mu$-neighborhood $U_{x}\subseteq X$ of $x$ such that $d\left[f_{ij}(y),F_{ij}(y)\right]<\varepsilon\mathtt{J}(y,z)$ for every $y\in U_{x}\cap\overline{E_{ij}}$.
Let ${s}:{U}\rightarrow{E_{ij}}$ be any function defined on a $\mu$-neighborhood of $x$ such that $\mathtt{J}\left[s(y),x\right]\leq L\mathtt{J}(y,x)$ near $x$. Therefore,
$\begin{split}d\left[f(y),F_{ij}(y)\right]&\leq d\left[f(y),f\left[s_{x}(y)\right]\right]+d\left[f_{ij}\left[s_{x}(y)\right],F_{ij}\left[s_{x}(y)\right]\right]\\ &\qquad+d\left[F_{ij}\left(y\right),F_{ij}\left[s_{x}(y)\right]\right]\\ &\leq d\left[f(y),f\left[s_{x}(y)\right]\right]+o\left(\mathtt{J}\left[s_{x}(y),x\right]\right)+K\mathtt{J}\left(s_{x}(y),y\right)\\ &=d\left[f(y),f\left[s_{x}(y)\right]\right]+o\left(\mathtt{J}\left(y,x\right)\right)\text{.}\end{split}$
We now use lemma 2.10 to find an appropriate function $s$. Let for every $y\in E_{ij}\cap B_{1/j}(x)$
$R(y)=\min\left\{{\frac{1}{3},\mathtt{J}(y,x)}\right\}$
$\tilde{B}_{y}=B_{y}\cap B_{R(y)\mathtt{J}(y,x)}(y)\text{.}$
Clearly each $\tilde{B}_{y}$ is $\left[\delta,R(y)\mathtt{J}(y,x)\right]$-full at $y$ and $R(y)$ depends only on $\mathtt{J}(y,x)$, therefore assumptions at point 2. in lemma 2.10 are satisfied and there exists ${s}:{U}\rightarrow{E_{ij}}$ such that $y\in\tilde{B}_{s(y)}$ for $y$ near $x$. In particular, $y\in B_{s(y)}$ then
$d[f(y),f[s(y)]]\leq i\mathtt{J}\left[y,s(y)\right]\text{.}$
Moreover, since $R(y)<1/2$, we also have $\mathtt{J}\left[y,s(y)\right]=o\left(\mathtt{J}(y,x)\right)$ therefore
$d[f(y),f[s(y)]]=o(\mathtt{J}(y,x))$
and $f$ is approximately $\mathcal{D}$-differentiable at $x$.
The converse if trivial. Indeed if every function ${f}:{X}\rightarrow{M}$ is approximately $\mathcal{D}$-differentiable a.e. on $A_{f}^{\mu}$ then to prove that $\mathcal{D}$ is complete we need only to accordingly extend any Lipschitz function ${g}:{C}\rightarrow{M}$ on all $X$ with $C$ closed. If $\tilde{g}$ is any extension of $g$ on all $X$ then we always have
${C}^{1}\cap A^{\mu}_{g}\subseteq A_{\tilde{g}}^{\mu}$
therefore $g$ is automatically a.e. approximately $\mathcal{D}$-differentiable, then $\mathcal{D}$ is complete.
Theorem 2.12.
Let $\mathcal{D}$ be a differentiably complete $X$-distribution for $M$, then every function ${f}:{A}\rightarrow{M}$ is a.e. $\mathcal{D}$-differentiable on $A_{f}$.
Conversely let $\mathcal{D}$ be an $X$-distribution such that every function ${f}:{X}\rightarrow{M}$ is $\mathcal{D}$-differentiable on $A_{f}$, then $\mathcal{D}$ is differentiably complete.
The first part of this result can be proved as in the previous theorem by setting $B_{y}=B_{3/j}(y)$ when $y\in E_{ij}$. Moreover, in the last passage we do not need lemma 2.10 to find $s$ but we can use directly proposition 2.3 such that $s$ is now defined in a neighborhood of $x$ instead of a $\mu$-neighborhood.
Now assume that every function ${f}:{X}\rightarrow{M}$ is $\mathcal{D}$-differentiable a.e. on $A_{f}$, we will prove that $\mathcal{D}$ is complete or equivalently that every Lipschitz function ${g}:{C}\rightarrow{M}$ defined on any closed subset $C\subseteq X$ is $\mathcal{D}$-differentiable a.e. on $C$. We want to define an extension of $g$ to whole $X$, an even though we cannot expect that such extension is Lipschitz it has useful properties:
Proposition 2.13.
If ${f}:{C}\rightarrow{M}$ is a generic function defined on a nonempty closed subset $C$ of $X$ then there exists ${\tilde{f}}:{X}\rightarrow{M}$ such that $\left.{\tilde{f}}\right|_{C}=f$ and $A_{f}\subseteq A_{\tilde{f}}$.
Proof.
We define $\tilde{f}$ in the following way: if $x\in C$ then we set $\tilde{f}(x)=f(x)$, instead if $x\notin C$ then we set $\tilde{f}(x)=f(y)$ where $y\in C$ is any point that satisfies
$\mathtt{J}\left(x,y\right)<2\operatorname{dist}(x,C)\text{.}$
Clearly $\tilde{f}$ is an extension of $f$, we need only to prove that $A_{f}\subseteq A_{\tilde{f}}$. Let $x\in A_{f}\cap\partial C$, therefore there exist $K,\delta>0$ such that if $y\in C\cap B_{\delta}(x)$ then $d[f(x),f(y)]\leq K\mathtt{J}\left(x,y\right)$.
Let $z\in B_{\delta/2}(x)\setminus C$, we can then find $y\in C\cap B_{\delta}(x)$ such that $\mathtt{J}\left(y,z\right)<2\operatorname{dist}(z,C)$ and $\tilde{f}(z)=f(y)$. Therefore,
$d[\tilde{f}(x),\tilde{f}(z)]=d[f(x),f(y)]\leq K\mathtt{J}\left(x,y\right)\leq K\left(\mathtt{J}\left(x,z\right)+2\operatorname{dist}(z,C)\right)\\ \leq 3K\mathtt{J}\left(x,z\right)$
which implies $x\in A_{\tilde{f}}$ with $\left\lvert\nabla\tilde{f}\right\rvert^{+}(x)\leq 3\left\lvert\nabla f\right\rvert^{+}(x)$. ∎
Now let ${f}:{C}\rightarrow{M}$ be Lipschitz and $C\subseteq X$ closed, then we have an extension ${\tilde{f}}:{X}\rightarrow{M}$ such that $C\subseteq A_{\tilde{f}}$. Since $\tilde{f}$ is $\mathcal{D}$-differentiable a.e. on $A_{\tilde{f}}$ we have also that $g$ is $\mathcal{D}$-differentiable a.e. on $C$, therefore $\mathcal{D}$ is complete.
3. Q-valued functions
On a generic metric space $(M,d)$ we can define the space of $Q$-points $\mathcal{A}_{Q}\left(M\right)$, where $Q$ is a positive integer, as the set of all measures $\mathbf{P}$ on $M$ that can be represented as the sum of $Q$ Dirac measures $\left\llbracket P_{1}\right\rrbracket+\left\llbracket P_{2}\right\rrbracket+\dotsb+\left\llbracket P_{Q}\right\rrbracket$ not necessarily distinct. Then a $Q$-valued function from a set $A$ to $M$ is just a function defined on $A$ with values on $\mathcal{A}_{Q}\left(M\right)$, clearly $1$-valued functions from $A$ to $M$ are equivalent to classical (single-valued) functions from $A$ to $M$.
Moreover, $Q$ single valued functions $f_{1},f_{2},\dotsc,f_{Q}$ can be used to compose a single $Q$-valued function $f(x)=\left\llbracket f_{1}(x)\right\rrbracket+\left\llbracket f_{2}(x)\right\rrbracket+\dotsb+\left\llbracket f_{Q}(x)\right\rrbracket$, converselt every $Q$-valued function $f$ can be globally decomposed into $Q$ single valued functions by choosing a total ordering on $M$. However, such decomposition is never useful except when $M$ is a really good space (usually $M={\mathbb{R}}$) because many useful properties of $f$ will not be inherited by such general decompositions.
On $\mathcal{A}_{Q}\left(M\right)$ we define the following Wassenstein metric:
$\mathcal{G}\left[\mathbf{A},\mathbf{B}\right]=\mathcal{G}\left[\sum^{Q}_{i=1}\left\llbracket A_{i}\right\rrbracket,\sum^{Q}_{i=1}\left\llbracket B_{i}\right\rrbracket\right]=\min_{\sigma\in\mathcal{S}_{Q}}\sqrt{\sum^{Q}_{i=1}d\left(A_{1},B_{\sigma(i)}\right)^{2}}$
where $\mathcal{S}_{Q}$ is the set of all the permutations on $\left\{{1,2,\dotsc,Q}\right\}$. Therefore a $Q$-valued function $f$ from a topological/metric/measurable space $\Omega$ to $M$ is continuous/Lipschitz/measurable if and only if ${f}:{\Omega}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is continuous/Lipschitz/measurable. Every measurable $Q$-valued function can be globally decomposed into $Q$ measurable single-valued functions (see [3]), however continuous $Q$-valued functions cannot usually be decomposed into $Q$ continuous single-valued functions, for example the following function
${z\in{\mathbb{C}}}\rightarrow{\left\llbracket z^{3/2}\right\rrbracket+\left\llbracket-z^{3/2}\right\rrbracket\in\mathcal{A}_{2}\left({\mathbb{C}}\right)}$
cannot be decomposed into two continuous functions in any neighborhood of the origin.
If $\mathcal{D}$ is an $X$-distribution for a complete metric space $M$ then we are able to derive another $X$-distribution $\mathcal{D}^{Q}$ but on $\mathcal{A}_{Q}\left(M\right)$: for every $\mathbf{P}=\sum^{Q}_{i=1}\left\llbracket P_{i}\right\rrbracket$ we say $\left[\mathbf{F}\right]\in\mathcal{D}^{Q}_{x_{0},\mathbf{P}}$ if and only if there exist $\left[F_{i}\right]\in\mathcal{D}_{x_{0},P_{i}}$ for every $i=1,2,\dotsc,Q$ such that
$\mathbf{F}(x)=\sum^{Q}_{i=1}\left\llbracket F_{i}(x)\right\rrbracket\ \forall x\text{ in a Borel neighborhood of }x_{0}$ (3.1)
if $\text{if }P_{i}=P_{j}\text{ then }\left[F_{i}\right]=\left[F_{j}\right]\text{.}$ (3.2)
Notice that $\mathcal{A}_{Q}\left(M\right)$ is always a complete metric space whenever $M$ is complete, and can be proved by induction on $Q$. This is clearly trivial for $Q=1$, thus assume both $M$ and $\mathcal{A}_{Q}\left(M\right)$ complete. Let $\mathbf{P}_{l}\in\mathcal{A}_{Q+1}\left(M\right)$ be a Cauchy sequence, by induction on $l$ we can easily find a subsequence in the form $\mathbf{P}_{h_{l}}=\left\llbracket P_{l}\right\rrbracket+\mathbf{P}^{\prime}_{l}$, where both $P_{l}\in M$ and $\mathbf{P}^{\prime}_{l}\in\mathcal{A}_{Q}\left(M\right)$ are Cauchy sequences that converge to some $P\in M$ and $\mathbf{P}^{\prime}\in\mathcal{A}_{Q}\left(M\right)$ respectively. Therefore $\mathbf{P}_{h_{l}}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ which in turn implies $\mathbf{P}_{l}\to\left\llbracket P\right\rrbracket+\mathbf{P}^{\prime}$ because $\mathbf{P}_{l}$ is Cauchy.
Many authors (see for example [3]) imposes condition (3.2) in their definition of differentiable $Q$-valued functions since it makes easier to work componentwise. For example if $M={\mathbb{R}}^{n}$ and $\left[\mathbf{F}\right]\in\mathcal{D}^{Q}$ is a differential germ for $f=\sum^{Q}_{i=1}\left\llbracket f_{i}\right\rrbracket$ at some $x\in{\mathbb{R}}^{k}$ with $f_{i}$ continuous at $x$ then we can always take the $i$-th component $F_{i}\in\mathcal{D}$ of $\mathbf{F}$ so that $\left[F_{i}\right]$ is a $\mathcal{D}$-differential for $f_{i}$ at $x$.
Instead other authors (for example [7]) do not use (3.2) in their definition of differentiable $Q$-valued functions, since with (3.2) we could have a sequence of $C^{1}$ $Q$-valued functions $f_{n}$ such that both $f_{n}$ and $\nabla f_{n}$ uniformly converge to some functions $f$ and $F$ respectively but $f$ is not differentiable everywhere! Therefore when one need to work with sequences of $C^{1}$ $Q$-valued functions it may be useful to require only (3.1) without (3.2).
However, if one need only a.e. differentiability (for example when dealing with Sobolev functions) then (3.2) is unimportant as we will see later. To show this, in this place we will always assume that germs in $\mathcal{D}^{Q}$ satisfy both (3.1) and (3.2).
Theorem 3.1.
For every $\mathbf{P}\in\mathcal{A}_{Q}\left(M\right)\setminus\left\{{Q\left\llbracket P\right\rrbracket}\mathrel{}\middle|\mathrel{}{P\in M}\right\}$ there exist $1\leq R<Q$, an open neighborhood $U$ of $\mathbf{P}$ and two continuous maps ${\pi_{1}^{\mathbf{P}}}:{U}\rightarrow{\mathcal{A}_{R}\left(M\right)}$, ${\pi_{2}^{\mathbf{P}}}:{U}\rightarrow{\mathcal{A}_{Q-R}\left(M\right)}$ such that
  • $\mathbf{Q}=\pi_{1}^{\mathbf{P}}(\mathbf{Q})+\pi_{2}^{\mathbf{P}}(\mathbf{Q})$ for every $\mathbf{Q}\in U$;
  • $\mathcal{G}\left[\mathbf{Q},\mathbf{R}\right]^{2}=\mathcal{G}\left[\pi_{1}^{\mathbf{P}}(\mathbf{Q}),\pi_{1}^{\mathbf{P}}(\mathbf{R})\right]^{2}+\mathcal{G}\left[\pi_{2}^{\mathbf{P}}(\mathbf{Q}),\pi_{2}^{\mathbf{P}}(\mathbf{R})\right]^{2}$ for every $\mathbf{Q},\mathbf{R}\in U$.
Proof.
Clearly there exist $1\leq R<Q$, $P$, $P_{R+1},P_{R+2},\dotsc,P_{Q}\in M$, and $\varepsilon>0$ such that $\mathbf{P}=R\left\llbracket P\right\rrbracket+\sum^{Q}_{i=R+1}\left\llbracket P_{i}\right\rrbracket$ and $d\left(P,P_{j}\right)\geq\varepsilon$ for every $R+1\leq j\leq Q$. Therefore, the subset $U\subseteq\mathcal{A}_{Q}\left(M\right)$ defined so that $\mathbf{A}\in U$ if and only if there exists $\sigma\in S_{Q}$ such that $d\left(P,A_{\sigma(i)}\right)<\varepsilon/3$ for every $i=1,2,\dotsc,R$ and $d\left(P_{i},A_{\sigma(i)}\right)<\varepsilon/3$ for every $i=R+1,R+2,\dotsc,Q$ is open and contains $\mathbf{P}$.
If $\mathbf{A}\in U$ we can assume without loss of generality that $\sigma(i)=i$, then we always have
$A_{1},A_{2},\dotsc,A_{R}\in B_{\varepsilon/3}(P)$
$A_{R+1},A_{R+2},\dotsc,A_{Q}\notin B_{2\varepsilon/3}(P)$
$\mathcal{G}\left[\mathbf{P},\mathbf{A}\right]^{2}=\sum^{R}_{i=1}d\left(P,A_{i}\right)^{2}+\sum^{Q}_{i=R+1}d\left(P_{i},A_{\eta(i)}\right)^{2}$
where $\eta\in S_{Q}$ with $\eta(i)=i$ for every $1\leq i\leq R$. We will prove only the last assertion. First of all assume that
$\mathcal{G}\left[\mathbf{P},\mathbf{A}\right]^{2}=\sum^{R}_{i=1}d\left(P,A_{\eta(i)}\right)^{2}+\sum^{Q}_{i=R+1}d\left(P_{i},A_{\eta(i)}\right)^{2}$
and set
$E=\left\{{i=1,2,\dotsc,R}\mathrel{}\middle|\mathrel{}{\eta(i)\geq R+1}\right\}$
$F=\left\{{i=R+1,R+2,\dotsc,Q}\mathrel{}\middle|\mathrel{}{\eta(i)\leq R}\right\}\text{.}$
Set $P_{i}=P$ for $i=1,\dotsc,R$ we clearly have $d\left(P_{i},A_{\eta(i)}\right)\geq 2\varepsilon/3$ if $i\in E\cup F$. Moreover, both $E$ and $F$ must have the same cardinality thus if $E$ is not empty then we can find another $\eta^{\prime}\in S_{Q}$ such that $\eta^{\prime}(i)=\eta(i)$ if $i\notin E\cup F$ and $d\left(P_{i},A_{\eta^{\prime}(i)}\right)<\varepsilon/3$ for every $i$. But this is impossible by our assumption on $\eta$.
Therefore we can set
${\pi_{1}^{\mathbf{P}}}:{\mathbf{A}\in U}\rightarrow{\sum^{R}_{i=1}\left\llbracket A_{i}\right\rrbracket\in\mathcal{A}_{R}\left(M\right)}$
${\pi_{2}^{\mathbf{P}}}:{\mathbf{A}\in U}\rightarrow{\sum^{Q}_{i=R+1}\left\llbracket A_{i}\right\rrbracket\in\mathcal{A}_{Q-R}\left(M\right)}\text{.}\qed$
Corollary 3.2.
Let $\Omega\subseteq X$ be open nonempty and ${f}:{\Omega}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ be continuous at $x\in\Omega\setminus\Lambda$ where
$\Lambda=\left\{{x\in\Omega}\mathrel{}\middle|\mathrel{}{f(x)=Q\left\llbracket P\right\rrbracket\text{ for some }P\in M}\right\}\text{.}$
Then there exist an open neighborhood $U$ of $x$ and ${g}:{U}\rightarrow{\mathcal{A}_{R}\left(M\right)}$, ${h}:{U}\rightarrow{\mathcal{A}_{S}\left(M\right)}$ with $Q=R+S$, $\left.{f}\right|_{U}=\left\llbracket g\right\rrbracket+\left\llbracket h\right\rrbracket$ and
$\left\lvert\nabla g\right\rvert^{+}(x),\left\lvert\nabla h\right\rvert^{+}(x)\leq\left\lvert\nabla f\right\rvert^{+}(x)\text{.}$
Theorem 3.3.
If an approximate $X$-distribution $\mathcal{D}$ for $M$ is complete then the $X$-distribution $\mathcal{D}^{Q}$ for $\mathcal{A}_{Q}\left(M\right)$ satisfying both (3.1) and (3.2) is complete too.
We use induction on $Q$. If $Q=1$ then we can use assumption (RC) on $M$, thus we can assume that theorem 3.3 holds for any $1\leq Q^{\prime}<Q$ in order to prove it for $Q$. Let ${f}:{C}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ be a Lipschitz function defined on a closed subset $C$ of $X$ and let $x\in C$.
If $x\notin\Lambda$ then by corollary 3.2 we have $\left.{f}\right|_{V}=\left\llbracket g\right\rrbracket+\left\llbracket h\right\rrbracket$ with $x\in A_{g}\cap A_{h}=V$ for some open neighborhood $V$ of $x$, thus we can use induction on $g$ and $h$ to prove that $f$ is a.e. differentiable on $V$. Since $C$ is separable we can cover $C\setminus\Lambda$ by countably many of such $V$, thus we immediately get that $f$ is a.e. differentiable on $C\setminus\Lambda$.
On $\Lambda$ we have instead $f(x)=Q\left\llbracket f_{a}(x)\right\rrbracket$ with ${f_{a}}:{\Lambda}\rightarrow{M}$ Lipschitz, therefore by induction hypothesis $f_{a}$ is a.e. approximately $\mathcal{D}$-differentiable on $A^{\mu}_{f_{a}}$. If $L\subseteq A^{\mu}_{f_{a}}$ is the set of points on which $f_{a}$ is approximately $\mathcal{D}$-differentiable with differential $[F]\in\mathcal{D}_{x,f_{a}(x)}$ then we can assume that $L\subseteq{\Lambda}^{1}\cap A_{f}^{\mu}$. In this way $F$ is approximately $\mathcal{D}^{Q}$-differentiable on every point of $L$ because
$\mathcal{G}\left[f(y),Q\left\llbracket F(y)\right\rrbracket\right]=\sqrt{Q}\ d\left[f_{a}(y),F(y)\right]$
for every $y\in\Lambda$.
Combining theorem 3.3 with theorem 2.12 we finally get the following generalized version of Stepanov’s theorem for $Q$-valued functions which generalizes theorem 1.2:
Theorem 3.4.
Let $(X,\mathtt{J},\mu)$ be a doubling metric measure space, $(M,d)$ be a complete metric space and $\mathcal{D}$ a complete (approximate) $X$-distribution on $M$. Then every $Q$-valued function ${f}:{X}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is a.e. (approximately) $\mathcal{D}^{Q}$-differentiable on $A_{f}$ (or $A_{f}^{\mu}$).
In particular, these properties hold whenever $M={\mathbb{R}}^{k}$ and $X$ is ${\mathbb{R}}^{n}$ or a $p$-Poincaré doubling metric measure space thanks to [2].